OpenStudy (anonymous):

is the equation -2(4-x)=2x-8 an identity

jimthompson5910 (jim_thompson5910):

what do you get when you distribute -2(4-x)

OpenStudy (anonymous):

i dont know :(

jimthompson5910 (jim_thompson5910):

hint: a*(b+c) = a*b + a*c

jimthompson5910 (jim_thompson5910):

that's the distributive property in general

OpenStudy (anonymous):

can you draw it

OpenStudy (anonymous):

-8 - 2x thats what i got

jimthompson5910 (jim_thompson5910):

-2 times (-x) = +2x or just 2x

jimthompson5910 (jim_thompson5910):

so you should have -8 + 2x

jimthompson5910 (jim_thompson5910):

that means -2(4-x)=2x-8 turns into -8 + 2x = 2x - 8

OpenStudy (anonymous):

ok how u get the positive sign

jimthompson5910 (jim_thompson5910):

the negative on the -2 and the negative on the -x they multiply to a positive

jimthompson5910 (jim_thompson5910):

(negative) times (negative) = positive

OpenStudy (anonymous):

ok but now what to do

jimthompson5910 (jim_thompson5910):

well since we can add two numbers in any order this means that -8 + 2x is really the same as 2x + (-8) but that turns into 2x - 8

jimthompson5910 (jim_thompson5910):

so -8 + 2x = 2x - 8 turns into 2x - 8

OpenStudy (anonymous):

how

jimthompson5910 (jim_thompson5910):

I just showed how

jimthompson5910 (jim_thompson5910):

you swap the terms in -8 + 2x to get 2x - 8

OpenStudy (anonymous):

ok thanks

jimthompson5910 (jim_thompson5910):

hopefully it's making sense?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

alright, that's great

OpenStudy (anonymous):

is 2x=3=7x a odentity or have no solution

OpenStudy (anonymous):

identity*

jimthompson5910 (jim_thompson5910):

there are too many equal signs

OpenStudy (anonymous):

2x + 3 =7x

jimthompson5910 (jim_thompson5910):

try to isolate x and tell me what you get

OpenStudy (anonymous):

3 = 5x

jimthompson5910 (jim_thompson5910):

keep going

OpenStudy (anonymous):

i dont know the rest

jimthompson5910 (jim_thompson5910):

you have 5x = 3 but you want something like x = ....

OpenStudy (anonymous):

please tell me i dont know

jimthompson5910 (jim_thompson5910):

5x means 5 times x

jimthompson5910 (jim_thompson5910):

what undoes multiplication?

OpenStudy (anonymous):

i dont know

jimthompson5910 (jim_thompson5910):

what operation is the opposite of multiplication?

OpenStudy (anonymous):

may you please understand i dont know understand none of this

jimthompson5910 (jim_thompson5910):

well something like subtraction is the opposite of addition so ________ is the opposite of multiplication

OpenStudy (anonymous):

divide

jimthompson5910 (jim_thompson5910):

good, so you divide both sides by 5 5x = 3 5x/5 = 3/5 x = 3/5

jimthompson5910 (jim_thompson5910):

so the solution to 2x + 3 =7x is x = 3/5 this is the only solution

OpenStudy (anonymous):

5(y-2) = -2(12-9y) +y